OpenStudy (anonymous):

The equation of the line that goes through the point (4,2) and is parallel to the line going through the points (−1,1) and (5,4) has what slope and equation?

OpenStudy (anonymous):

First do you know how to find the slope between the points (-1,1) and (5,4)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This is the desired slope for the equation that you are trying to find. Once you have a slope (m) and your point (here it would be (4,2)) you can plug these into y=mx+b and find your b component. For example say your slope is 4/5 and your point is (2,1) 1=(4/5)*2 +b ===> 1=8/5 +b ===> -3/5=b so your desired equation of the line with slope m=4/5 going through the point (2,1) is y=(4/5)x + (-3/5)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome. if you have any further questions I would be happy to help.